OpenStudy (anonymous):

subject: derivatives,... If dc/dx = 0, where c is a constant, then what is dx/dx ?

OpenStudy (anonymous):

\[\frac{ d }{dx }\left( c \right)=0\] \[\frac{ d }{ dx }\left( x \right)^{n}=nx ^{n-1}\] put n=1 and get the solution.

OpenStudy (anonymous):

Thanks.